Help!!!!SOS!!!!! I have been given this formula and given this instructions, I have been at it for 3 hours and I'm completely lost!!!!
Instructions X2 - 2X- 13 =0
a) Move the constant term to the right side of the equation.
b) Multiply each term in the equation by four times the coefficient of the x squared term.
c) Square the coefficient of the original x term and add it to both sides of the equation.
d) Take the square root of both sides.
e) Set the left side of the equation equal to the positive square root of the right side and solve for x.
f) Set the left side of the equation equal to the negative square root of the right side and solve for x.

Answer by richwmiller(17219)


There is no reason to do step b)
X^2-2X-13 = 0
X^2-2X=13
step c is wrong
square 1/2 of the coefficient of x
2/2=1 squared equals 1
(X-1)^2=14
x-1=+-sqrt(14)
X = 1+sqrt(14)
X = 1-sqrt(14)